Practice Exercise– Future Value and Present Value

                    ( worth 10 points extra credit) – due on Monday 4/13

  1. [ST 5-1 Problem on p.232]

Assume you have $10,000 to deposit in a bank for 3 years. Bank A compounds interest annually, Bank B compounds interest semi-annually , Bank C compounds interest quarterly . All three banks have a stated annual interest rate 4%

    (a) What amount would you have after three years, in each bank?

    (b)What APY (annual percentage yield) or EAR (effective annual interest rate) would you earn in each bank?

   (c) Based on your answers in (a) and (b), which bank should you deal with? Why?
Answer #1

a.

Bank A = Future Value = Deposit * (1 + Interest per Period)^Number of Periods

Bank A = Future Value = 10000 * 1.04^3 = $11248.64

Bank B = Future Value = Deposit * (1 + Interest per Period)^Number of Periods

Bank B = Future Value = 10000 * 1.02^6 = $11261.62

Bank C = Future Value = Deposit * (1 + Interest per Period)^Number of Periods

Bank C = Future Value = 10000 * 1.01^12 = $11268.25

b.

Computation of APY

Bank A = same as nominal interest rate = 4%

Bank B = (1 + Interest per Period)^Number of Periods in years] - 1 = (1 + 0.02)^2] - 1 = 4.04%

Bank C = (1 + Interest per Period)^Number of Periods in years] - 1 = (1 + 0.01)^4] - 1 = 4.06%

c. based on answers in a and b i would suggest to invest in Bank C which is offering quarterly compounding interest

Please Upvote if satisfied